OpenStudy (anonymous):

sneakers that regulary sell for $72 are being sold for 30% off. the sales tax is 8 %. what will the final cost , including tax for the sneakers

OpenStudy (therealmeeeee):

ok We're gonna tear this up. It's simple, but it will take a bit of patience, so buckle up. Ready? Let's go. You don't know the original price. You know the percent off. You know the sale price. We're in business. Let's hammer this thing. Here's how to work the problem.... We don't know the original price, but we know that a percentage of it has been deducted from it (that original price) to give us a sale price, okay? Some percent off the original price is the sale price. Here's the trick. Look at the percent off. Now look at 100% minus the percent off. This new percentage represents how much of the original cost the final cost is. Got it? Another way to say that is that our new (calculated) percentage times the original price equals the sale price. Make sense? Let's pick something easy and give it a test drive.

OpenStudy (therealmeeeee):

Hope I helped

OpenStudy (anonymous):

thnks

OpenStudy (therealmeeeee):

Welcome :)

OpenStudy (anonymous):

is ur name tyrek u look familar

OpenStudy (jdoe0001):

\(\bf \begin{array}{ccllll} quantity&percentage\\ \hline\\ 72&100\\ x&100-30 \end{array}\implies \cfrac{72}{x}=\cfrac{100}{100-30}\implies \square \\ \quad \\ \square \textit{ is then 70% of }\$72\textit{ now adding 8% of tax then} \\ \quad \\ {\color{red}{ 8}}\%\ of \ \square =\cfrac{{\color{red}{ 8}}}{100}\times \square \qquad thus \qquad \square +\left(\cfrac{{\color{red}{ 8}}}{100}\times \square\right)\textit{ is the total amount}\)